### v1

# Optomechanical Systems with Linear and Quadratic Position Couplings: Dynamics and Optimal Estimation

Abstract
We study the dynamics of an optomechanical system consisting of a single-mode optical field coupled to a mechanical oscillator, where the nonlinear interaction includes both linear and quadratic terms in the oscillator’s position. We present a full analytical solution to this quantum mechanical Hamiltonian problem by employing the formalism of two-phonon coherent states. Quantum estimation theory is applied to the resulting state of the optical field, with a focus on evaluating the classical and quantum Fisher information for estimating the strength of the quadratic coupling. Our estimation scheme considers both standard and balanced homodyne photodetection, assuming an initial optical state prepared as a superposition of vacuum and single-photon states. We show that balanced homodyne detection can saturate the quantum Fisher information, thus reaching the ultimate precision bound for estimating the quadratic coupling. Additionally, we investigate the effect of thermal noise on the quantum Fisher information in a realistic experimental context.
I Introduction
Optomechanical systems [1] are at the center of a technological revolution [2] that is based on the interaction between light and matter to study fundamental physics [3] and produce innovative devices, spanning from signal processors [4] to transductors [5] and quantum amplifiers [6, 7]. Achieving these goals requires a thorough understanding of the optomechanical system, which can be attained through precise measurements of its parameters. Numerous studies [8, 9, 10, 11, 12, 13] have focused on understanding how to estimate the parameters of an optical cavity coupled to a mechanical oscillator via radiation pressure.
In constructing effective estimation frameworks, the first essential step is to define a statistical model for the data [14, 15]. In quantum mechanics, this requirement translates to obtaining an exact solution for the time evolution of the quantum state. Formally, this yields a family of quantum states parameterized by the unknown variables to be estimated. However, due to the complexity introduced by the high-dimensional parameter spaces commonly encountered in optomechanical systems, it is often necessary to introduce simplifications to the physical model. These approximations facilitate the formulation of an minimal model that remains tractable while still capturing the essential physics. A detailed derivation of the optomechanical Hamiltonian is presented in the foundational works [16, 17]. Two principal approximations are typically employed to simplify this Hamiltonian. First, if the mechanical oscillator evolves adiabatically slowly compared to the frequency separation between optical modes, intermodal photon scattering can be neglected. Under this condition, the system can be effectively described using a single optical cavity mode. Second, in the linear approximation, the radiation–mechanical interaction is simplified by expanding the coupling Hamiltonian to first order in the mechanical position operator. The first objective of this paper is to construct a statistical model by deriving the exact time-evolving quantum state governed by the optomechanical Hamiltonian, under the adiabatic approximation while extending the standard linear interaction to include second-order (quadratic) terms in the mechanical displacement.
The Hamiltonian under investigation has been previously discussed in comprehensive reviews [1] and has found applications in various contexts, including quantum nondemolition measurements of the phonon number of the mechanical mode [18], as well as in optomechanically induced parametric oscillations [19]. Mathematical treatments often focus on the corresponding Heisenberg equations of motion, their expectation values, and employ approximations to derive a finite set of coupled differential equations for the system dynamics [20]. Our focus is on obtaining an exact solution for the time-evolving quantum state, which, to our knowledge, has not been previously derived. We employ the formalism of two-photon coherent states [21], adapted here to describe the mechanical oscillator, allowing us to call them as two-phonon coherent states. By working in the Fock basis of the single-mode optical field, the Hamiltonian becomes block diagonal, with each block corresponding to a fixed photon number. The dynamics within each block are then governed by the evolution of the photon number-related two-phonon coherent states. The result describes the joint quantum state of the optical field and the mechanical oscillator. Since, in practice, measurements are typically performed on the optical field, we trace out the mechanical degrees of freedom. We therefore focus on the resulting reduced state of the optical field. This reduced optical state serves as the basis for the quantum estimation procedure.
Here, we adopt a frequentist approach and focus on the lower bounds of the variance for any unbiased estimator. Even without explicitly constructing the estimators, these bounds serve as benchmark values for assessing the performance of those used in experiments. In this context, we analyze the quantum Fisher information (QFI), which represents the lower bound in the Cramér Rao inequality [22, 23], for the optical field state as a function of the unknown strength of the optomechanical coupling, which depends quadratically on the mechanical position operator. This is compared with the classical Fisher information (CFI), obtained for standard and balanced-homodyne photodetection measurement schemes. The QFI serves as a measure of the sensitivity of the optical state to changes in the quadratic optomechanical coupling constant; in other words, it quantifies how much information the quantum state contains about this parameter. On the other hand, the CFI assesses the potential to evaluate this sensitivity using a set of classical measurements. For simplicity, we assume that the mechanical oscillator is initially in a thermal state, while the optical field is prepared in a superposition of the vacuum and single-photon Fock states. It is important to note that the cavity is assumed to be lossless, and its quantum state can be accessed without inducing any disturbance.
This paper is organized as follows. In Sec. II, we describe the analytical model and we solve the equations of motion, with explicit dependence upon the linear and quadratic (the mechanical position operator) optomechanical coupling constants. In Sec. III we introduce the quantum and classical Fisher information and the projective measures describing the standard and the balanced-homodyne photodetection schemes. In Sec. IV we show the dynamics of the optical state affected by quadratic optomechanical coupling, and we evaluate the quantum and classical Fisher information, while analyzing the performance of the two measurement schemes. Finally, in Sec. V we draw our conclusions and outlooks.
II Model
We consider a single-mode Fabry-Pérot optical cavity with a suspended dielectric membrane that is free to oscillate. The electromagnetic field inside the cavity exerts radiation pressure on the mechanical oscillator and causes it to oscillate around its equilibrium position. The Hamiltonian of the optomechanical system can be written by summing the contribution from the optical and the mechanical subsystems as [17]
$\hat{H}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\Omega^{2}\hat{x}^{2}+\hbar\omega(\hat{x})\hat{a}^{\dagger}\hat{a},$ (1)
where $\hat{x}$ is the position operator of the mechanical system, $\hat{p}$ is its momentum operator, $m$ is its mass, and $\Omega$ is the mechanical frequency of motion. The operators $\hat{a},\hat{a}^{\dagger}$ are the photonic ladder operators with the commutation relation $[\hat{a},\hat{a}^{\dagger}]=1$, and $\omega(x)$ is the resonant optical frequency of the cavity, which depends on the position $x$ of the mechanical oscillator. If we expand the optical frequency $\omega(x)$ around the equilibrium value $x_{0}$, we obtain
$\omega(x)$$\approx$ $\omega(x_{0})+\omega^{\prime}(x_{0})(x-x_{0})+\frac{\omega^{\prime\prime}(x_{0})}{2!}(x-x_{0})^{2}$ (2)
$=$ $\omega_{c}+g_{1}x+g_{2}x^{2},$
where $\omega^{(n)}(x_{0})$ denotes the $n$th derivative of $\omega(x)$ evaluated at the point $x_{0}$. Here, $g_{1}$ is the linear, and $g_{2}$ is the quadratic optomechanical coupling constant. The parameter $\omega_{c}$ is identified as the frequency of the single mode of the radiation field. By considering the position as an operator, we can write the optomechanical Hamiltonian as
$\hat{H}^{\text{quad}}=\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\Omega^{2}\hat{x}^{2}+\hbar\omega_{c}\hat{a}^{\dagger}\hat{a}+\hbar g_{1}\hat{a}^{\dagger}\hat{a}\hat{x}+\hbar g_{2}\hat{a}^{\dagger}\hat{a}\hat{x}^{2}.$ (3)
It is worth mentioning that the single-mode approach requires the motion of the mechanical oscillator to be adiabatically slow, so the model can ignore the scattering of photons from the mode to other cavity modes [16].
In order to solve the dynamics, we notice that the Hamiltonian (3) is block diagonal in the photonic Fock basis $|n\rangle_{c}$ ($n\in{\mathbb{N}}_{0}$), i.e. it can be written as
$\hat{H}^{\text{quad}}=\sum_{n}\hat{H}_{n},\quad\hat{H}_{n}=\hat{P}_{n}\hat{H}^{\text{quad}}\hat{P}_{n},$ (4)
where $\hat{P}_{n}=|n\rangle_{c}\langle n|$ is the projector on the subspace with $n$ photons. The Hamiltonian $\hat{H}_{n}$ reads
$\hat{H}_{n}=\hbar n\omega_{c}\hat{I}+\frac{\hat{p}^{2}}{2m}+\frac{1}{2}m\bigg{(}\Omega^{2}+\frac{2\hbar g_{2}n}{m}\bigg{)}\hat{x}^{2}+\hbar g_{1}n\hat{x},$ (5)
where $\hat{I}$ is the identity operator on the Hilbert space of the mechanical oscillator. By setting $\Omega_{n}=\sqrt{\Omega^{2}+\frac{2\hbar g_{2}n}{m}}$, we can define the ladder mechanical operators $\hat{b}_{n},\hat{b}^{\dagger}_{n}$ as
$\hat{b}_{n}$ $=$ $\sqrt{\frac{m\Omega_{n}}{2\hbar}}\bigg{(}\hat{x}+i\frac{\hat{p}}{m\Omega_{n}}\bigg{)},$ (6)
$\hat{b}^{\dagger}_{n}$ $=$ $\sqrt{\frac{m\Omega_{n}}{2\hbar}}\bigg{(}\hat{x}-i\frac{\hat{p}}{m\Omega_{n}}\bigg{)}.$ (7)
We stress that the ladder operators related to different indices $\hat{b}_{n},\hat{b}_{l}$ with $n\neq l$ do not commute. In fact, the relations between them are given by the following Bogoliubov-like transformation rules:
$\hat{b}_{l}$ $=$ $\frac{1}{2}\sqrt{\frac{\Omega_{l}}{\Omega_{n}}}\bigg{[}\bigg{(}1-\frac{\Omega_{n}}{\Omega_{l}}\bigg{)}\hat{b}_{n}^{\dagger}+\bigg{(}1+\frac{\Omega_{n}}{\Omega_{l}}\bigg{)}\hat{b}_{n}\bigg{]},$
$\hat{b}_{l}^{\dagger}$ $=$ $\frac{1}{2}\sqrt{\frac{\Omega_{l}}{\Omega_{n}}}\bigg{[}\bigg{(}1-\frac{\Omega_{n}}{\Omega_{l}}\bigg{)}\hat{b}_{n}+\bigg{(}1+\frac{\Omega_{n}}{\Omega_{l}}\bigg{)}\hat{b}_{n}^{\dagger}\bigg{]}.$ (8)
This must be taken into account when solving the dynamics of the system, because any initial state of the mechanical oscillator is represented by the Fock basis $|m\rangle_{0}$ ($m\in{\mathbb{N}}_{0}$) defined by the ladder operators $\hat{b}_{0},\hat{b}^{\dagger}_{0}$. The Hamiltonian (5) assumes the form
$\hat{H}_{n}=\hbar\omega_{n}\hat{I}+\hbar\Omega_{n}\hat{b}^{\dagger}_{n}\hat{b}_{n}+\hbar g_{1,n}(\hat{b}_{n}+\hat{b}^{\dagger}_{n}),$ (9)
where $\omega_{n}=n\omega_{c}+\Omega_{n}/2$ and $g_{1,n}=ng_{1}\sqrt{\frac{\hbar}{2m\Omega_{n}}}$. By applying the Baker-Campbell-Hausdorff formula, we get [24]
$e^{-i\hat{H}_{n}t/\hbar}=e^{-i\Phi_{n}(t)}e^{\eta_{n}(t)\hat{b}_{n}^{\dagger}-\eta_{n}^{*}(t)\hat{b}_{n}}e^{-i\Omega_{n}\hat{b}^{\dagger}_{n}\hat{b}_{n}t},$ (10)
where we have introduced the functions
$\Phi_{n}(t)$ $=$ $\omega_{n}t-\frac{g^{2}_{1,n}}{\Omega_{n}^{2}}[\Omega_{n}t-\sin(\Omega_{n}t)],$ (11)
$\eta_{n}(t)$ $=$ $\frac{g_{1,n}}{\Omega_{n}}(e^{-i\Omega_{n}t}-1).$ (12)
Given an initial state $\rho(0)$, the time evolution of the system is given by the von Neumann equation
$\hat{\rho}(t)=e^{-i\hat{H}^{\text{quad}}t/\hbar}\hat{\rho}(0)e^{i\hat{H}^{\text{quad}}t/\hbar}.$ (13)
We are interested in cases where no initial correlations between the single-mode optical field and the mechanical oscillator are considered. Therefore, we choose an initial state of the form
$\hat{\rho}(0)=\hat{\sigma}(0)\otimes\mathfrak{\hat{m}}(0),$ (14)
where $\hat{\sigma}(0)$ and $\mathfrak{m}$ are the optical and the mechanical density operator, respectively.
Now, we consider the density operator of the mechanical oscillator in the following representation:
$\mathfrak{\hat{m}}(0)=\int d^{2}\alpha\,P(\alpha,\alpha^{*})|\alpha\rangle_{0}\langle\alpha|.$ (15)
where $d^{2}\alpha=d\mathrm{Re}\{\alpha\}\,d\mathrm{Im}\{\alpha\}$, $\lvert{\alpha}\rangle_{0}$ is a coherent state, and $P(\alpha,\alpha^{*})$ is a Glauber-Sudarshan quasidistribution [25, 26]. Due to the fact that $\hat{H}^{\text{quad}}$ is block diagonal in the photonic Fock basis, we choose
$\hat{\sigma}(0)=\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert,$ (16)
where the matrix entries $a_{n,m}$ are subject to the conditions
$\mathrm{Tr}\{\hat{\sigma}(0)\}=1\quad\text{and}\quad\hat{\sigma}(0)\geq 0,$ (17)
i.e., $\hat{\sigma}(0)$ is a positive semidefinite operator with trace one.
To evaluate the time evolution of the state, we employ the Bogoliubov transformations (II) to calculate the action of $e^{-iH_{n}t/\hbar}$ on the eigenstates of the number operator $\hat{N}_{n}=\hat{b}^{\dagger}_{n}\hat{b}_{n}$. They define a Fock basis $|m\rangle_{n}$ ($m\in{\mathbb{N}}_{0}$), whose properties have been extensively investigated in the literature [21, 27, 28, 29]. The displacement operator $\hat{D}_{n}(\alpha)=e^{\alpha\hat{b}^{\dagger}_{n}-\alpha^{*}\hat{b}_{n}}$ has the following property
$\hat{D}_{n}(\alpha)|0\rangle_{n}=|\alpha\rangle_{n},$ (18)
where $\lvert{\alpha}\rangle_{n}$ is an example of a two-phonon coherent state. We note that the concept of two-photon coherent states was originally developed for quantized electromagnetic fields  [21]; however, in this work, we apply the formalism to an oscillator with mass. With the help of the unitary squeeze operator
$\hat{S}(z)=e^{\frac{1}{2}(z^{*}\hat{b}^{2}_{0}-z\hat{b}^{\dagger 2}_{0})},\quad z=r\,e^{i\theta},$ (19)
we have the transformation
$\lvert{m}\rangle_{n}$ $=$ $\hat{S}(z_{n})\lvert{m}\rangle_{0},\quad z_{n}=-r_{n}\,e^{i\theta_{n}},$ (20)
$\cosh r_{n}$ $=$ $\frac{1}{2}\sqrt{\frac{\Omega_{n}}{\Omega}}\bigg{(}1+\frac{\Omega}{\Omega_{n}}\bigg{)},\quad\theta_{n}=0.$ (21)
Now, we can evaluate Eq. (13) as
$\hat{\rho}(t)$ $=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\hat{H}_{n}t/\hbar}\hat{\mathfrak{m}}(0)e^{i\hat{H}_{m}t/\hbar}$ (22)
$=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}\int d^{2}\alpha\,P(\alpha,\alpha^{*})D_{n}\left[\eta_{n}(t)\right]e^{-i\Omega_{n}\hat{N}_{n}t}|\alpha\rangle_{0}\langle\alpha|e^{i\Omega_{m}\hat{N}_{m}t}D^{\dagger}_{m}\left[\eta_{m}(t)\right].$
In the next step, we introduce two identity operators
$\hat{I}=\frac{1}{\pi}\int\lvert{\beta}\rangle_{n}\langle{\beta}\rvert\,d^{2}\beta=\frac{1}{\pi}\int\lvert{\gamma}\rangle_{m}\langle{\gamma}\rvert\,d^{2}\gamma$ (23)
to get
$\hat{\rho}(t)$ $=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}$
$\times$ $\int\frac{P(\alpha,\alpha^{*})}{\pi^{2}}\hat{D}_{n}\left[\eta_{n}(t)\right]e^{-i\Omega_{n}\hat{b}^{\dagger}_{n}\hat{b}_{n}t}\lvert{\beta}\rangle_{n}\langle\beta\lvert{\alpha}\rangle_{0}\langle{\alpha}\rvert\gamma\rangle_{m}\langle{\gamma}\rvert e^{i\Omega_{m}\hat{b}^{\dagger}_{m}\hat{b}_{m}t}\hat{D}^{\dagger}_{m}\left[\eta_{m}(t)\right]\,d^{2}\alpha\,d^{2}\beta\,d^{2}\gamma$
$=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}\int\frac{P(\alpha,\alpha^{*})}{\pi^{2}}\lvert{\eta_{n}(t)+\beta e^{-i\Omega_{n}t}}\rangle_{n}\langle\beta\lvert{\alpha}\rangle_{0}\langle{\alpha}\rvert\gamma\rangle_{m}\langle{\eta_{m}(t)+\gamma e^{-i\Omega_{m}t}}\rvert$
$\times$ $e^{\frac{1}{2}\left[\eta_{n}(t)\beta^{*}e^{i\Omega_{n}t}-\eta^{*}_{n}(t)\beta e^{-i\Omega_{n}t}\right]}e^{\frac{1}{2}\left[\eta_{m}^{*}(t)\gamma e^{-i\Omega_{m}t}-\eta_{m}(t)\gamma^{*}e^{i\Omega_{m}t}\right]}\,d^{2}\alpha\,d^{2}\beta\,d^{2}\gamma,$ (25)
where we have used the relations:
$e^{-i\Omega\hat{b}^{\dagger}_{n}\hat{b}_{n}t}|\alpha\rangle_{n}$ $=$ $|\alpha e^{-i\Omega t}\rangle_{n},$ (26)
$\hat{D}_{n}(\beta)|\alpha\rangle_{n}$ $=$ $e^{\frac{1}{2}(\beta\alpha^{*}-\beta^{*}\alpha)}|\beta+\alpha\rangle_{n}.$ (27)
The remainder of the calculation will focus solely on the optical density operator, $\hat{\sigma}(t)$, at time $t$. This approach is justified by the fact that, in most practical scenarios, the optical field is the only component of the system that is experimentally accessible. The operator $\hat{\sigma}(t)$ is obtained by performing a partial trace over the mechanical degrees of freedom, and it can be expressed as
$\hat{\sigma}(t)=\mathrm{Tr}_{\text{mech}}\{\hat{\rho}(t)\}=\sum^{\infty}_{n,m=0}a_{n,m}(t)|n\rangle_{c}\langle m|,$ (28)
where
$a_{n,m}(t)=a_{n,m}e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}\int\frac{P(\alpha,\alpha^{*})}{\pi^{3}}{}_{n}\langle\beta\lvert{\alpha}\rangle_{0}\langle{\alpha}\rvert\gamma\rangle_{m}\,{}_{0}\langle\delta\lvert{\eta_{n}(t)+\beta e^{-i\Omega_{n}t}}\rangle_{n}\,{}_{m}\langle{\eta_{m}(t)+\gamma e^{-i\Omega_{m}t}}\rvert\delta\rangle_{0}\,d^{2}\alpha\,d^{2}\beta\,d^{2}\gamma\,d^{2}\delta$ (29)
To derive this equation and to subsequently make use of Eq. (LABEL:eq:A1), we have introduced the identity operator
$\hat{I}=\frac{1}{\pi}\int\lvert{\delta}\rangle_{0}\langle{\delta}\rvert\,d^{2}\delta.$ (30)
The integral can be evaluated explicitly under the assumption that the mechanical system initially is in a thermal state:
$P(\alpha,\alpha^{*})=\frac{e^{-|\alpha|^{2}/n_{\text{th}}}}{\pi n_{\text{th}}}$ (31)
with the average phonon number
$n_{\text{th}}=\biggl{[}\exp\biggl{(}\frac{\hbar\Omega}{k_{\text{B}}T}\biggr{)}-1\biggr{]}^{-1},$ (32)
where $k_{\text{B}}$ is the Boltzmann constant and $T$ the temperature of the initial thermal state. The expression for $a_{n,m}(t)$ reduces to a Gaussian integral, yielding
$a_{n,m}(t)$ $=$ $I\int_{\mathbb{R}_{8}}\exp(-\frac{1}{2}\textbf{x}^{T}A\textbf{x}+b^{T}\textbf{x}+c)d^{8}\textbf{x}$ (33)
$=$ $I\sqrt{\frac{(2\pi)^{8}}{\textbf{det}(A)}}\exp(\frac{1}{2}b^{T}A^{-1}b+c),$
where $I$ reads
$I=\frac{a_{n,m}e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}}{\pi^{4}n_{\text{th}}\frac{\Omega_{n}+\Omega}{2\sqrt{\Omega\Omega_{n}}}\frac{\Omega_{m}+\Omega}{2\sqrt{\Omega\Omega_{m}}}}.$ (34)
The complete expressions for $A$, $b$ and $c$ are provided in Appendix B, all derived from the overlaps of two-phonon coherent states [29], as explicitly given in Eq. (LABEL:eq:A1). The coefficients $a_{n,m}(t)$ of the optical density operator provide a complete description of the interaction between the single mode of the radiation field and the single vibrational mode of the mechanical oscillator, under the assumption that all losses and sources of decoherence are neglected. We now use these results to estimate the strength of the quadratic optomechanical coupling.
III Quantum and Classical Fisher information
In an estimation problem, the goal is to infer the value of an unknown parameter by inspecting datasets coming from the measurement. In the literature, various approaches to this problem are presented, which broadly fall into two main categories: the Bayesian and frequentist frameworks [14, 15]. In this work, we adopt the frequentist perspective, wherein the parameter is assumed to have a true, fixed (i.e., nonrandom) value. The central question then becomes: what is this true value, and what is the optimal strategy to estimate it? In quantum metrology, this question is often addressed by comparing the classical Fisher information (CFI) with the quantum Fisher information (QFI).
Given a certain measurement procedure, the CFI quantifies the information that can be extracted from a given state and it is upper-bounded by the QFI, which is the maximization of the CFI over all quantum measurements. Thus, QFI is a property of the state, whereas CFI is a property of both the state and the measurement procedure. Their importance comes from the Cramér-Rao theorem [14] and its quantum mechanical version [22], which states the relation between the Fisher information and the minimum attainable variance of any unbiased estimator. It reads
$\text{Var}_{g}\geq\frac{1}{N_{\text{meas}}\text{CFI}}\geq\frac{1}{N_{\text{meas}}\text{QFI}}.$ (35)
While the construction of estimators will not be addressed in this work, it plays a crucial role in estimation theory and can be particularly challenging when dealing with probability distributions arising from quantum systems [30]. To be more specific, the measurements that we perform on the system return a set of data, which are then fed to an estimator, a function of the data whose outcome is an estimate of the unknown parameter. Whether the variance of our estimate reaches the minimum given by the inverse of the CFI, following the Cramér-Rao theorem (35), depends on the estimator itself. For example, the maximum likelihood estimator is able to saturate the lower bound given by the CFI for asymptotically large data sizes [14, 15]. We point out to the reader that there exist other, less well known versions of the Cramér-Rao theorem for biased estimators [8], or when adopting the Bayesian approach [31].
In quantum mechanics, the measurement procedure is described by a Positive Operator-Valued Measure (POVM), a collection of operators $\{\Pi_{i}\}_{i}$, such that any $\Pi_{i}\geq 0$ and $\sum_{i}\Pi_{i}=\mathds{1}$. The CFI of the state $\rho$ and the measurement $\Pi$ can be calculated as
$\text{CFI}=\sum_{i}P_{i}[g]\bigg{(}\frac{\partial\ln P_{i}[g]}{\partial g}\bigg{)}^{2},$ (36)
where $P_{i}[g]=\text{Tr}[\rho_{g}\Pi_{i}]$ is the probability of getting the outcome $i$ when the true value of the parameter is $g$.
The QFI could be obtained by maximizing Eq. (36) over all the possible POVMs, but this is generally a mission impossible. Fortunately, there is a simple expression of it [22]:
$\text{QFI}=\text{Tr}[\rho_{g}L_{g}^{2}]$ (37)
which is written in terms of the symmetric logarithmic derivative (SLD) $L_{g}$ [32], which satisfies the relation
$\frac{\partial\rho_{g}}{\partial g}=\frac{1}{2}(L_{g}\rho+\rho L_{g}).$ (38)
The SLD can be found by solving the Lyapunov equation (38), with the solution
[…]
We mention here that when the density operator has the peculiar property that $\sigma^{2}=\sigma-h\mathbf{I}$, with $h\in\mathbb{R}$, which always happens when $\sigma$ describes a two-level system, the SLD assumes the simplified form [33]
$L_{g}=2\partial_{\theta}\rho-\frac{1}{2}\partial_{\theta}\mathcal{P}\rho^{-1},$ (40)
where $\mathcal{P}=\text{Tr}[\rho^{2}]$ is the purity of the system and is related to $h$ by
$\mathcal{P}=2h+1,$
which can be used to ease the computation of the QFI. In the following we are going to apply Eq. (40) to calculate the QFI for the case when the cavity is populated by a superposition of zero-photon and one-photon states.
In the case of a $2\times 2$ system we can define an analogue of the photodetection and the balanced homodyne detection (BHD) measurements in terms of their projective value measures (PVM). For the photodetection, the corresponding PVM is $\{\Pi_{\text{ph}}\}$, with
$\Pi_{\text{ph}}(0)=|0\rangle\langle 0|$ (41)
$\Pi_{\text{ph}}(1)=|1\rangle\langle 1|.$
The BHD [34] enables the measurement of the observable
$\hat{X}_{\phi}=\frac{\hat{a}e^{i\phi}+\hat{a}^{\dagger}e^{-i\phi}}{2}.$ (42)
which, we rewrite to account for the two-level system scenario that we are considering by substituting the operators $\hat{a},\hat{a}^{\dagger}$ with the $2\times 2$ counterparts $\sigma^{-},\sigma^{+}$ respectively. The corresponding PVM $\{\Pi_{X_{\phi}}\}$ is given by the projectors on the two eigenstates $|e_{0,1}\rangle$ of $X_{\phi}$,
$\Pi_{X_{\phi}}(k)=|e_{k}\rangle\langle e_{k}|,\quad k=1,2.$ (43)
IV Results
We start our analysis by considering a simple scenario in which the initial photonic state $\hat{\sigma}(0)$ can be written on the Hilbert space spanned by the zero-photon state $|0\rangle$ and the one-photon state $|1\rangle$. Thus, $\hat{\sigma}(0)$ is a $2\times 2$ positive semi-definite density matrix that can be written as
$\hat{\sigma}(0)=\frac{1}{2}(\mathbf{I}+r_{x}\sigma_{x}+r_{y}\sigma_{y}+r_{z}\sigma_{z}),$ (44)
where $\sigma_{i}$ with $i=x,y,z$ are the Pauli matrices and the coefficients $r_{i}$ are real numbers s.t. $r_{x}^{2}+r_{y}^{2}+r_{z}^{2}\leq 1$, with the equality satisfied when $\hat{\sigma}(0)$ is a pure state.
IV.1 Unitless Model
To present the system dynamics in an intuitive manner, we first consider a unitless model in which all relevant physical parameters are set to one and unitless, while the linear coupling coefficient $g_{1}$ is set to zero and the quadratic $g_{2}$ to 0.01. See Table 1.
UM RS
[…]
$n_{th}$ $1$ $10^{5}$
[…]
$g_{2}$ $0.01$ $2\pi\times 4.46\times 10^{24}$ $[\text{Hz}/\text{m}^{2}]$
Table 1: Table of parameter values used in the unitless model (UM) and the real-world system (RS). SM values are unitless; RS values are in SI units.
In this section, we consider either the pure state
$|\psi_{s}\rangle=\sqrt{s}|0\rangle+\sqrt{1-s}|1\rangle,$ (45)
with $r_{x}(0)=2\sqrt{s(1-s)}$, $r_{y}(0)=0$ and $r_{z}=(2s-1)$, or the mixed state defined as
$\rho_{s}=\frac{1}{2}|\psi_{s}\rangle\langle\psi_{s}|+\frac{1}{2}|\psi_{1-s}\rangle\langle\psi_{1-s}|,$ (46)
both parametrized by the single parameter $s$. Note that the coefficients $r_{x}$ and $r_{y}$ of the pure states $|\psi_{s}\rangle$ and $|\psi_{1-s}\rangle$ coincide. In the following, we assign the value $s=0.2$, while a similar dynamics has been observed for different choices of $s.$
IV.1.1 System dynamics
Figure 1: The trajectory in the X-Y plane of the Bloch sphere of the evolution of the quantum states (45),(46) parameterized by $s=0.1$ (red) and $s=0.2$(blue). The system is evolved from time is 0 to $4\pi/\delta,$ which contains two complete cycles of the fast dynamic oscillation. The first cycles are denoted by solid lines and the second by dotted lines. To make the spiraling evolution more visible, the values of Bloch sphere coordinates $r_{x}$ and $r_{y}$ are scaled by $r^{2}=r_{x}^{2}+r_{y}^{2}.$ The evolution depends only on $s$, regardless on the purity of the state.
In Fig. 1 we show the parametric evolution of the optical state (33) expressed in terms of the coefficients $r_{x}$ and $r_{y}$, allowing the system evolve between $t=0$ and $t=\frac{4\pi}{\delta},$ with $\delta=\omega_{1}-\omega_{0}.$ The evolution is shown for two choices of the parameter $s$ of Eqs. (45) and (46). In fact, by solving the integral in Eq. (33), we can see that the diagonal elements of the optical density matrix are time-independent and therefore $r_{z}$ is a constant. Thus, the projection on the X-Y plane of the Bloch sphere suffices to describe the evolution of the optical system, while its value is independent from the purity $\mathcal{P}$ of the system, and can be parametrized solely by $s$. Note that, in order to make the features of the trajectories more visible, the axes have been scaled accordingly to $r_{x}/r^{2}$ and $r_{y}/r^{2}$, where $r^{2}=r_{x}^{2}+r_{y}^{2}.$ Comparing the first cycle in X-Y plane to the second, one observes the inward spiral of the coordinates towards the center of the Bloch sphere, which represents the maximally mixed state. Thus, the inward spiral shows a decrease in purity. Furthermore, we show in Figure 2 the evolution of the $r_{x}$ component. After about 50 fast cycles, the system spirals into the innermost point corresponding to the lowest purity, and then spirals outward again until it is back at the initial, outermost coordinates corresponding to the highest purity. We plot the time evolution of the purity $\mathcal{P}$ in Fig. 3 for a pure and mixed state with $s=0.2$.
Figure 2: Time evolution of $r_{x}$, the $x$ coordinate of the system in Bloch sphere, here plotted for the initial pure state $|\psi_{s}\rangle$ with $s=0.2$. Time is labeled by periods of complete fast oscillation, two of which are shown in Figure 1. Enveloping the fast oscillation is the beating oscillation $\Delta=\Omega_{1}-\Omega$, roughly $100$ times slower than the fast frequency.
Figure 3: The time evolution of the purity $\mathcal{P}$ for the pure (blue continuous line) and the mixed (red dashed line) states defined in Eqs. (45),(46) respectively, with $s=0.2$. Time is scaled by $\frac{2\pi}{\Delta}$, the natural oscillation period. In this figure, four oscillation periods have been plotted.
IV.1.2 Spectral Analysis
The time dependence of the optical coefficients as expressed in Eq. (33) is given by the term $\det(A)$, with main frequency being the beat frequency $\Delta=\Omega_{1}-\Omega$, and by the $\exp(i[\Phi_{1}(t)-\Phi_{0}(t)])$ factor, which contains the faster frequency $\delta=\omega_{1}-\omega_{0}$, while the last term $\exp(\frac{1}{2}b^{T}A^{-1}b+c)$ is a time-independent constant. In order to reveal the underlying frequencies, we performed a Fourier Transform on the dynamics of the $r_{x}$ coefficient,
[…]
which captures the off-diagonal time changing components of the optical density matrix. The two frequencies with larger amplitude are $\Delta$ and $\delta=\omega_{c}+\Delta/2$, while successive frequencies have an exponentially decaying amplitude, as shown in Figs. 4 and  5 respectively.
Figure 4: The absolute value of the Fourier transform of the coefficient $r_{x}(t)$, plotted in the range of the slow frequencies. The peaks are approximately multiples of $\frac{1}{3}\frac{\Delta}{2\pi}$.
Figure 5: The absolute value of the Fourier transform of the coefficient $r_{x}(t)$, plotted in the range of the fast frequencies. The first peaks is $\frac{\delta}{2\pi}$. This term originates from the $\exp(i[\Phi_{1}-\Phi_{0}]),$ whose exponent is proportional to $\omega_{1}-\omega_{0}$ in our system with $g_{1}=0.$ The peaks display an exponential decay in relative amplitude.
Looking at Fig. 4, we see that the major peaks are multiples of the $32\pi\Delta/$ 3. Conversely, the peaks in Fig. 5 represent the different periods of revolution of the $r_{x}$ coefficient, as can be also observed in Fig. 1, by comparing the position of the two $t=2\pi/\delta$ evolutions.
IV.1.3 Quantum Fisher Information
In Fig. 6 we show the QFI of the pure state $|\psi_{s}\rangle$ parameterized by $s=0.2$ numerically calculated using the definition (37). Note that a similar behavior is observed for other choices of the parameter $s$. The evolution of the QFI of the pure state presents sharp peaks at $t=2\pi/\Delta$. In Fig. 7, we show the temporal evolution of the QFI for the initial mixed state (46) with $s=0.2$. We calculated the QFI using the simplified analytic expression for $L_{g}$ given in Equation (40). Even for an initial mixed state, the periodicity of the peaks of the QFI coincide with the periodicity of the purity of the system. Around the peaks, the off-diagonal elements of the density matrix show greater sensitivity to changes of the $g_{2}$ parameter, as can be verified by solving Eq. (33). Contrary to the pure state, the peaks for the mixed state present a fast oscillation as we describe next.
Figure 6: Quantum Fisher Information as a function of time for the pure state with $s=0.2$, calculated with numerical integration. Periodic peaks are observed at the regular time interval of $2\pi/\Delta,$ entirely overlapping with the evolution of the purity.
Figure 7: Quantum Fisher Information as a function of time for the mixed state with $s=0.2$, calculated analytically. Periodic peaks are observed at the regular time interval of $2\pi/\Delta.$
IV.1.4 Classical Fisher Information
By solving Eq. (33), we see that the diagonal elements remain time-independent throughout the system evolution. Thus, the CFI calculated from the photodetection PVM, Eq. (41) is null, as information can be derived from the dynamics of the off-diagonal elements only. Conversely, the set of PVM related to the BHD, as defined in Eq. (43) provide information about the off-diagonal elements, and they are chosen to calculate the CFI. When the phase $\phi=0$ they reduce to
$\Pi_{0}=|+\rangle\langle+|,\;\Pi_{1}=|-\rangle\langle-|,$
where |±$|\pm\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)$, which lead to the measurement probability distribution:
$P_{0}=\frac{1}{2}(1+\rho_{01}+\rho_{10}),\;P_{1}=\frac{1}{2}(1-\rho_{01}-\rho_{10}).$
Note that $\frac{1}{2}(\rho_{01}+\rho_{10})$ is simply the real part of both the off-diagonal elements.
We compare the CFI obtained with the above described set of PVM to the QFI in Fig. 8, where we show the result around the first QFI peak set at $t\approx\Delta/2\pi$ in the case of an initial mixed state with $s=0.2$. The CFI is fast oscillating, with period $2\pi/\delta$ and its peak is slightly shifted with respect to the QFI peak. However, we find that when the CFI is maximum, its value corresponds to the QFI value. Similar behavior is obtained both for different choices of $s$ and for an initial pure state.
Figure 8: Classical Fisher Information (black dashed line) and Quantum Fisher Information (red continuous line) as a function of time zoomed in around the first period of the system’s beating oscillation, i.e., $t=\frac{2\pi}{\Delta}$.
IV.2 Real-world system
We take the numerical values from the paper [35], which describes a suspended SiN membrane that oscillates with frequency $\Omega=2\pi\times 134\text{ kHz}$ placed in the middle of a Fabry-Pérot optical cavity of length $L=67\text{ mm}$ driven by an external laser of wavelength $\lambda=1064\text{ nm}$. This configuration ensures a low effective mass of the harmonic oscillator, $m\approx 50\text{ pg}$. The fundamental frequency of the cavity changes with the position $x$ of the membrane as
$\omega(x)=\frac{c}{L}\arccos\bigg{[}|r|\cos\frac{4\pi x}{\lambda}\bigg{]},$ (48)
with $r=0.42$, from which we can calculate the expansion expressed in Eq. (2). This setup guarantees a relatively high magnitude of the quadratic optomechanical coupling $g_{2}\approx 2\pi\times 24\text{ kHz}\cdot\text{nm}^{-2}$ obtained by placing the membrane either in a node, or in an antinode of the magnetic field, when $g_{1}\approx 0$. Furthermore, in Refs.  [36, 37] it is shown how, by tilting the membrane of a small angle, we can induce an interaction between different transverse modes of the electromagnetic field, leading to an increased coupling constant $g_{2}\approx 2\pi\times 4.46\text{ MHz}\cdot\text{nm}^{-2}$. We consider the system to be at room temperature $T=300K$, and we consider the initial state of the mechanical system, (15) to be a thermal state with $n_{\text{th}}\approx 10^{5}$, Eq. (32). These values are reported in Table 1.
Using these values, we found that the first peak of the QFI is $\approx 5.5\times 10^{15}$ for an initial mixed state $\hat{\rho}_{s}$ with $s=0.2$. We analyze how the choice of the parameter $\phi$ influences the CFI. While the behavior shown in Fig. 8 is reflected even with the real-world values, we see that a different choice of $\phi$ can lead the CFI to saturate the QFI. In Fig. 9 we plot at $t=2\pi(\frac{1}{\Delta}+\frac{1}{\delta})$, the CFI as a function of the phase $\phi.$
Figure 9: QFI (red dashed line) and CFI (black continuous line) for the real-world system, calculated at $t=2\pi(\frac{1}{\Delta}+\frac{1}{\delta})$ as a function of the quadrature phase $\phi$.
Finally, we investigate the influence of temperature on the QFI. In Fig. 10 we plot the QFI as a function of the temperature for the different peaks at $t=2\pi N/\Delta,$ with $N=1,10,$ and $100$
Figure 10: Quantum Fisher Information as a function of the system temperature plotted at $t=2\pi N/\Delta$, with $N=1,10,$ and $100$. Depending on the temperature the largest peak of the QFI is located at different periods of oscillation.
When the temperature is close to zero the state of the mechanical system is substantially the ground state of the mechanical oscillator. This means that the optical system is interacting with a very weak field and consequently the information about the optomechanical coupling is low. At the opposite, when the temperature is large, the mechanical system is close to the maximally mixed state, which is persistent, regardless of the value of $g_{2}.$ Even in this case, the information retrieval about $g_{2}$ is small. In between, there is a temperature where the QFI is at its maximum.
From Fig. 10 we see that the largest peak of the QFI is located at a different period of oscillation that depends on the temperature of the mechanical system. This behavior is opposite for cold and hot temperatures. For low $T,$ although the mechanical system is approximately in the ground state, the information grows at each cycle, as the information available to the optical system cumulates. For high $T,$ the mechanical bath tends to destroy all the available information and the peaks of the QFI decline at each cycle.
V Conclusion and perspectives
In this work we studied the dynamics of an optomechanical system with quadratic coupling in the position operator. We found that the each photon number brings a separate contribution to the global Hamiltonian, thus allowing us to isolate these contributors and solve the temporal evolution of the density operator. Under the assumption that the optical and mechanical quantum states are factorized at initial time, and considering the mechanical system in a thermal state, we were able to derive the full description of the optical density matrix.
We studied the evolution on the Bloch sphere of a simple system with the initial optical state being a superposition of zero and one-photon states. The optical system has a purity that oscillates with a period frequency given by the beat frequency between mechanical system associated with different photons number.
Hence, we analyzed the sensitivity of the optical field to changes of the quadratic optomechanical coupling by calculating the quantum Fisher information on the reduced density matrix. The effect of the quadratic optomechanical coupling becomes evident when the system regains its original purity, as the off-diagonal elements of the optical density matrix store the information about the coupling constant. We studied the classical Fisher information associated with the measurement of the quadrature operator, finding that it is able to saturate the quantum Fisher information when the phase of the quadrature and the phase aligns with the phase of the off-diagonal terms.
Furthermore, we studied the dependence of the quantum Fisher information to the temperature of the mechanical system, showing that it reaches its maximum at an intermediate temperature that allows the state of the mechanical system to be highly populated while being far from the maximally mixed state.
Our study describes the procedure to gain the most information on the optomechanical coupling constant out of the optical field inside the cavity. Future works need to assess whether our result remains valid when the measurement is performed on the output field, which can be obtained by filtering the input-output relation to the detector frequency [38, 9].
Acknowledgements.
J.Z.B. acknowledges support from AIDAS-AI, Data Analytics and Scalable Simulation, which is a Joint Virtual Laboratory gathering the Forschungszentrum Jülich and the French Alternative Energies and Atomic Energy Commission, and from the Hungarian National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004).
Appendix A Relevant properties of the two-phonon coherent state
In this appendix we write the overlap between the two-phonon coherent state $|\beta\rangle_{n}$ and the coherent state $|\alpha\rangle.$ It is [21]
$\langle\alpha|\beta\rangle_{n}$ $=$ $\frac{1}{\sqrt{\mu_{n}}}e^{-\frac{1}{2}(|\alpha|^{2}+|\beta|^{2}+\frac{\nu_{n}}{\mu_{n}}\alpha^{*2}-\frac{\nu_{n}^{*}}{\mu_{n}}\beta^{2}-\frac{2}{\mu_{n}}\alpha^{*}\beta)};$
where the parameters $\mu_{n},\nu_{n}$ are taken from the Bogoliubov transformations
$\hat{b}_{n}$ $=$ $(\mu_{n}\hat{b}+\nu_{n}\hat{b}^{\dagger})$ (50)
$\hat{b}_{n}^{\dagger}$ $=$ $(\mu_{n}^{*}\hat{b}^{\dagger}+\nu_{n}^{*}\hat{b}).$ (51)
and satisfy the relation
$|\mu_{n}|^{2}-|\nu_{n}|^{2}=1.$ (52)
Their value is obtained by comparing the Bogoliubov equations (II), with the above formula, yielding:
$\mu_{n}=\frac{\Omega_{n}+\Omega}{2\sqrt{\Omega\Omega_{n}}},\quad\nu_{n}=\frac{\Omega_{n}-\Omega}{2\sqrt{\Omega\Omega_{n}}}.$ (53)
With the proper substitutions we can calculate all the implicit terms in Eq. (33).
Appendix B Expressions for the multivariate Gaussian integral
In this appendix we explicitly write the elements of Eq. (33). Consistently with the expression found in Eq. (29) and introducing the identity as in Eq. (30), we define the x vector as
$\textbf{x}=\Big{(}\mathrm{Re}\{\alpha\},\mathrm{Im}\{\alpha\},\mathrm{Re}\{\beta\},\mathrm{Im}\{\beta\},\mathrm{Re}\{\delta\},\mathrm{Im}\{\delta\},\mathrm{Re}\{\gamma\},\mathrm{Im}\{\gamma\}\Big{)},$ (54)
and the matrix $A$ as
[…]
with $M$ given by
$\begin{pmatrix}\dfrac{2}{N_{th}}+F_{+}&iF_{-}&-\frac{1}{\mu_{n}}&\frac{i}{\mu_{n}}&0&0&-\frac{1}{\mu_{m}}&-\frac{i}{\mu_{m}}\\ iF_{-}&\dfrac{2}{N_{th}}-F_{+}&-\tfrac{i}{\mu_{n}}&-\tfrac{1}{\mu_{n}}&0&0&\tfrac{i}{\mu_{m}}&-\tfrac{1}{\mu_{m}}\\ -\tfrac{1}{\mu_{n}}&-\tfrac{i}{\mu_{n}}&-\dfrac{\nu_{n}\bigl{(}1+e^{-2i\Omega_{n}t}\bigr{)}}{\mu_{n}}&i\,\dfrac{\nu_{n}\bigl{(}1-e^{-2i\Omega_{n}t}\bigr{)}}{\mu_{n}}&-\tfrac{e^{-i\Omega_{n}t}}{\mu_{n}}&\tfrac{ie^{-i\Omega_{n}t}}{\mu_{n}}&0&0\\ \tfrac{i}{\mu_{n}}&-\tfrac{1}{\mu_{n}}&i\,\dfrac{\nu_{n}\bigl{(}1-e^{-2i\Omega_{n}t}\bigr{)}}{\mu_{n}}&\dfrac{\nu_{n}\bigl{(}1+e^{-2i\Omega_{n}t}\bigr{)}}{\mu_{n}}&-\tfrac{ie^{-i\Omega_{n}t}}{\mu_{n}}&-\tfrac{e^{-i\Omega_{n}t}}{\mu_{n}}&0&0\\ 0&0&-\tfrac{e^{-i\Omega_{n}t}}{\mu_{n}}&-\tfrac{ie^{-i\Omega_{n}t}}{\mu_{n}}&F_{+}&-iF_{-}&-\tfrac{e^{\,i\Omega_{m}t}}{\mu_{m}}&\tfrac{ie^{\,i\Omega_{m}t}}{\mu_{m}}\\ 0&0&\tfrac{ie^{-i\Omega_{n}t}}{\mu_{n}}&-\tfrac{e^{-i\Omega_{n}t}}{\mu_{n}}&-iF_{-}&-F_{+}&-\tfrac{ie^{\,i\Omega_{m}t}}{\mu_{m}}&-\tfrac{e^{\,i\Omega_{m}t}}{\mu_{m}}\\ -\tfrac{1}{\mu_{m}}&\tfrac{i}{\mu_{m}}&0&0&-\tfrac{e^{\,i\Omega_{m}t}}{\mu_{m}}&-\tfrac{ie^{\,i\Omega_{m}t}}{\mu_{m}}&-\dfrac{\nu_{m}\bigl{(}1+e^{2i\Omega_{m}t}\bigr{)}}{\mu_{m}}&i\,\dfrac{\nu_{m}\bigl{(}e^{2i\Omega_{m}t}-1\bigr{)}}{\mu_{m}}\\ -\tfrac{i}{\mu_{m}}&-\tfrac{1}{\mu_{m}}&0&0&\tfrac{ie^{\,i\Omega_{m}t}}{\mu_{m}}&-\tfrac{e^{\,i\Omega_{m}t}}{\mu_{m}}&i\,\dfrac{\nu_{m}\bigl{(}e^{2i\Omega_{m}t}-1\bigr{)}}{\mu_{m}}&\dfrac{\nu_{m}\bigl{(}1+e^{2i\Omega_{m}t}\bigr{)}}{\mu_{m}}\end{pmatrix},$
where the following shorthands were used:
$F_{+}=\tfrac{\nu_{n}}{\mu_{n}}+\tfrac{\nu_{m}}{\mu_{m}},\;F_{-}=\tfrac{\nu_{n}}{\mu_{n}}-\tfrac{\nu_{m}}{\mu_{m}}.$
The vector $B$ has the form
$B=-\frac{1}{2}\begin{pmatrix}0\\ 0\\ 2\Bigl{\{}\Re\left[\eta_{n}(t)\right]\cos(\Omega_{n}t)-\Im\left[\eta_{n}(t)\right]\sin(\Omega_{n}t)-\eta_{n}(t)\,\tfrac{\nu_{n}}{\mu_{n}}e^{-i\Omega_{n}t}\Bigr{\}}-\eta_{n}(t)\,e^{i\Omega_{n}t}+\eta^{*}_{n}(t)\,e^{-i\Omega_{n}t}\\ 2\Bigl{\{}\Re\left[\eta_{n}(t)\right]\sin(\Omega_{n}t)+\Im\left[\eta_{n}(t)\right]\cos(\Omega_{n}t)-i\,\eta_{n}(t)\,\tfrac{\nu_{n}}{\mu_{n}}e^{-i\Omega_{n}t}\Bigr{\}}+i\,\eta_{n}(t)\,e^{i\Omega_{n}t}+i\,\eta^{*}_{n}(t)\,e^{-i\Omega_{n}t}\\ -\tfrac{2}{\mu_{n}}\,\eta_{n}(t)-\tfrac{2}{\mu_{m}}\,\eta^{*}_{m}(t)\\ \tfrac{2i}{\mu_{n}}\,\eta_{n}(t)-\tfrac{2i}{\mu_{m}}\,\eta^{*}_{m}(t)\Bigr{)}\\ 2\Bigl{\{}\Re\left[\eta_{m}(t)\right]\cos(\Omega_{m}t)-\Im\left[\eta_{m}(t)\right]\sin(\Omega_{m}t)-\eta^{*}_{m}(t)\,\tfrac{\nu_{m}}{\mu_{m}}e^{i\Omega_{m}t}\Bigr{\}}-\eta^{*}_{m}(t)\,e^{-i\Omega_{m}t}+\eta_{m}(t)\,e^{i\Omega_{m}t}\\ 2\Bigl{\{}\Re\left[\eta_{m}(t)\right]\sin(\Omega_{m}t)+\Im\left[\eta_{m}(t)\right]\cos(\Omega_{m}t)+i\,\eta^{*}_{m}(t)\,\frac{\nu_{m}}{\mu_{m}}e^{i\Omega_{m}t}\Bigr{\}}-i\,\beta^{*}_{m}(t)\,e^{-i\Omega_{m}t}-i\,\eta_{m}(t)\,e^{i\Omega_{m}t}\end{pmatrix}.$ (55)
Finally, the scalar $c$ has the following explicit form:
$c\;=\;-\tfrac{1}{2}\Bigl{[}\lvert\eta_{n}(t)\rvert^{2}+\lvert\eta_{m}(t)\rvert^{2}\;-\;\frac{\nu_{m}}{\mu_{m}}\,\eta^{*2}_{m}(t)\;-\;\frac{\nu_{n}}{\mu_{n}}\,\eta_{n}^{2}(t)\Bigr{]}.$